How much energy (J) is required to heat 24.1 g of water from 23.6 oC to 33.39 oC?
1
Expert's answer
2019-10-17T07:42:52-0400

Initial temperature of water"(T_1)=23.6\degree C"

Final temperature of water"(T_2)=33.39\degree C"

Specific heat capacity of water"=4.184 J"

Heat required for the change in temperature"=mC(T_2-T_1)=24.1\times 4.184\times (33.39-23.6)=987.169J"